two bodies of masses 2kg &8kg have kinetic energies 16J&24J respectively. when they are moving in the same direction . they collide each other. the total momentum after collision is
1
Expert's answer
2012-08-17T10:09:35-0400
KE = m * V^2 / 2, so V = sqrt(2 * KE / m).
The total momentum P of the two
bodies is the same before the
collision as it is after the collision.
P =
m1 * V1 + m2 * V2
P = m1 * sqrt(2 * KE1 / m1) + m2 * sqrt(2 * KE2 / m2)
P
= sqrt(2 * KE1 * m1) + sqrt(2 * KE2 * m2)
P = sqrt(2 * 16 * 2) + sqrt(2 * 24
* 8) = 27.6 kg*m/s.